if the distance between the earth and the sun become 4 times ,then the time period becomes?

Grade:9

1 Answers

Arun
6 years ago
 

The ratio of the time periods would be {new is say T1, actual is say T2}

 

 

 

Note that T is directly proportional to √(R3)

 

 

 

T1/T2 = √([R*4]3)/√(R3) = 8

 

 

 

Hence T1 = T2* 8= 365 * 8

 

 

 

Which is approx 2920 days or 8 years